A scuba diver sees light reflected from the water’s surface. At what angle relative to the water’s surface will this light be completely polarized?

answer to 3 decimal places


1
Expert's answer
2021-06-02T13:06:45-0400

Answer:-

(note- external source like internet help is taken to get the exact values )

it is given that the refracting medium is water,

we can find the refractive index refracting medium , thus we have

"n_2=1.333"

and "n_1 (air)=1.000293"

and we know that

"tan(\theta)=\frac{n_2}{n_1}\\
\theta=tan^{-1}(\frac{n_2}{n_1})"

putting the values we get

"\theta=tan^{-1}(\frac{1.33}{1.000293})\\
\approx\boxed{53.115^o} \ answer"